OpenStudy (anonymous):

find the derivative of [sin^4(a^2x)]

terenzreignz (terenzreignz):

ahh.. chain rule Like a function hit-list :D Your expression, let's express it like \[\huge \left[ \sin \left( a^{2x} \right) \right]^{4}\] Key to chain rule is to identify the 'outermost' function. What is it?

terenzreignz (terenzreignz):

by outermost, like in this example, it's the \[x^{4}\] So how would you differentiate that?

terenzreignz (terenzreignz):

Yes, that's right. So now, we "killed" ;) the outermost function. \[\huge 4\left[ \sin \left( a^{2x} \right) \right]^{3}\] So now move on to the next level... Which is \[\huge \sin (a^{2x})\] What's the outermost function now?

terenzreignz (terenzreignz):

ok... Well, it doesn't matter what it was for now, let's work with \[\huge 4\left[ \sin \left( a^{2}x \right) \right]^{3}\] Does that look better now? :)

terenzreignz (terenzreignz):

No... for many many reasons... patience :D

terenzreignz (terenzreignz):

We got rid of the outermost function \[x^{4}\] And we got, for now:\[4\left[ \sin \left( a^{2}x \right) \right]^{3}\]Leave that to one side, we'll get back to it later. Now deal with the next 'outermost' function. Which happens to be sin x. How would you differentiate it?

terenzreignz (terenzreignz):

whoa there... differentiate sin x for now...

terenzreignz (terenzreignz):

You only touch the outermost functions, slowly working your way into the innermost.

terenzreignz (terenzreignz):

I have a question... did you forget that a is a constant? ;)

OpenStudy (anonymous):

Haha, I was not aware

terenzreignz (terenzreignz):

You did say d/dx so it should not affect a (this is not entirely true, but accept it for now)

terenzreignz (terenzreignz):

And lol, be patient differentiate the "sin" first before that thing inside the sin function :P

terenzreignz (terenzreignz):

And when I say differentiate the "sin" first, I mean leave everything else untouched...

OpenStudy (anonymous):

so cos(a^2 x)?

terenzreignz (terenzreignz):

Yes, so you "add that to your hitlist" ;) lol, I just like to think of things that way... \[\huge 4\left[ \sin \left( a^{2}x \right) \right]^{3}\left[ \cos \left( a^{2}x \right) \right]\] But we're not done... After the sin, there is another function inside \[a^{2}x\] how would you differentiate this?

OpenStudy (anonymous):

is that just 2a?

terenzreignz (terenzreignz):

2a? again, a^2 is just a constant...

OpenStudy (anonymous):

Haha sorry. Well, I still can't grasp this xD

terenzreignz (terenzreignz):

Well, what don't you get?

OpenStudy (anonymous):

just the a^2x

OpenStudy (anonymous):

I guess Ive already drilled in my head that its supposed to be 2a

terenzreignz (terenzreignz):

if it was d/da, then yes, that is what you should do but it's d/dx, as you said...

OpenStudy (anonymous):

is just 1?

terenzreignz (terenzreignz):

What is just 1?

OpenStudy (anonymous):

well in relation to x so a^2 (1)

OpenStudy (anonymous):

or just a^2

terenzreignz (terenzreignz):

I'm afraid I don't get you...

OpenStudy (anonymous):

Haha its fine, Ive always had issues with chain rule, I dont know why

terenzreignz (terenzreignz):

Better fix that.... chain rule is used extensively :)

OpenStudy (anonymous):

I know :(

OpenStudy (anonymous):

Well, what would I do for? 4[sin(a^2 (x))]3[cos(a^2(x))]I

terenzreignz (terenzreignz):

Well... So far, you have \[\huge 4\left[ \sin \left( a^{2}x \right) \right]^{3}\left[ \cos \left( a^{2}x \right) \right]\] All that's left is for you to multiply that with the derivative (WITH RESPECT TO X) of \[\huge a^{2}x\] which is just...?

OpenStudy (anonymous):

I feel like its such a trick question.. Hmmm..

terenzreignz (terenzreignz):

Imagine angels whispering in your ear in heavenly voices... "a is a constant, a is a constant, a is a constant...."

OpenStudy (anonymous):

Well its not 1, 2a, nor a^2

terenzreignz (terenzreignz):

Why not a^2 ? If it were 9x, what would be its derivative?

OpenStudy (anonymous):

9?

terenzreignz (terenzreignz):

That's right Now, what's the derivative of \[\huge 3^{2}x\] ?

OpenStudy (anonymous):

I mean 9

terenzreignz (terenzreignz):

Well, which is it?

OpenStudy (anonymous):

.. 6x?

terenzreignz (terenzreignz):

You... applied... power rule... to a constant... come on \[\huge 9x = 3^{2}x\] So why would they have different derivatives...?

OpenStudy (anonymous):

So by that statement it just remains a^2 x?

OpenStudy (anonymous):

considering 'a' is a constant

terenzreignz (terenzreignz):

Ok, let me just clear something up... The derivative of kx where k is a constant is k Regardless of how weird a constant k is. It can be as simple as c or as complicated as \[\frac{\ln c + \sin^{2}c}{\cos (c^{2} + 4c + 9)}\] It doesn't matter, as long as it's a CONSTANT. In a nutshell: THE DERIVATIVE OF A CONSTANT TIMES x IS THAT CONSTANT.

terenzreignz (terenzreignz):

What's a constant? For now, treat as a constant anything that does not involve x, or a function of x.

OpenStudy (anonymous):

Its a variable that always retains the same value no matter what, correct?

terenzreignz (terenzreignz):

The mere fact that it's a constant means it's not a variable...

OpenStudy (anonymous):

Haha I'm sorry, I understand that, just my words got mixed up :#

terenzreignz (terenzreignz):

Right, so knowing that, is the derivative of \[a^{2}x\] clear to you now?

OpenStudy (anonymous):

So if its just the constant, a^2 x = a?

OpenStudy (anonymous):

If not, I'll accept the fact I'm a complete derp

terenzreignz (terenzreignz):

Ok, I said that the derivative of kx is k. basically, I just dropped the x. Please do just that, even in the case of \[a^{2}x\]

OpenStudy (anonymous):

so its just a^2?

terenzreignz (terenzreignz):

YES

OpenStudy (anonymous):

derpderpderp

terenzreignz (terenzreignz):

Well, do you understand now?

OpenStudy (anonymous):

no, do I put that in the sin or cos part?

OpenStudy (anonymous):

now*

terenzreignz (terenzreignz):

no, you've already dealt with that by differentiating the sin into cos earlier, that's why you go further in...

terenzreignz (terenzreignz):

I meant, first you differentiated the ^4 part And then you differentiated the sin part finally, you differentiate the a^2 x part

OpenStudy (anonymous):

And thats just a^2

terenzreignz (terenzreignz):

Yes, and you multiply that to \[\huge 4\left[ \sin \left( a^{2}x \right) \right]^{3}\left[ \cos \left( a^{2}x \right) \right]\] And you'll be done.

OpenStudy (anonymous):

Thank you very much :) Haha, and thanks for putting up with my.. ignorance

terenzreignz (terenzreignz):

No problem. :)